A stock has an expected return of 11.6%. Given a risk-free rate of 3.3% and a market return of 14%, what is the β of this stock? Enter your answer rounded to 2 DECIMAL PLACES.

β=   
Answer #1

Expected return = Rf + β(Rm - Rf)

0.116 = 0.033 + β(0.14 - 0.033)

0.116 = 0.033 + β(0.107)

β(0.107) = 0.116 - 0.033

β(0.107) = 0.083

β = 0.083 / 0.107

β = 0.78